# SCORE: Story Coherence and Retrieval Enhancement for AI Narratives

### v1

@@ 5 Conclusion
We present SCORE, a novel LLM-based framework designed to address the critical challenges of long-term coherence and emotional consistency in AI-generated narratives. By integrating a hybrid architecture—combining Dynamic State Tracking, Context-Aware Summarization, and Hybrid Retrieval (TF-IDF + semantic embeddings)—SCORE leverages Retrieval-Augmented Generation (RAG) to enforce contextual alignment across episodic storytelling. Experimental validation on multi-genre datasets demonstrates SCORE’s superiority over baseline LLMs, with marked improvements in coherence, emotional consistency, and reduction in hallucinations, while maintaining robustness against noisy inputs. The framework’s modular design supports incremental knowledge graph construction and multi-LLM compatibility, offering a scalable solution for industrial applications. However, limitations persist, including reliance on retrieval accuracy for key-item continuity and computational overhead from hierarchical summarization. Future work will explore lightweight memory compression and real-time sentiment alignment to further optimize performance.
Appendix A Theoretical Analysis
A.1 Proof of the Continuity Analysis of Critical Item State Correction
@@
$P\big{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1},\dots,S(t_{0})=s_{0}\big{)}=P\big{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1}\big{)}.$ (7)
@@
Thus, no transition from destroyed or lost back to active or any non‐terminal state is allowed.
@@
$\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}=-\ln\Bigl{(}P\bigl{(}S_{k+1}=\text{active}\big{|}S_{k}=\text{destroyed}\bigr{)}\Bigr{)}>0,$ (13)
@@
$E=\sum_{j\neq k}\Delta E\big{(}S_{j},S_{j+1}\big{)}+\Delta E\big{(}\text{destroyed},\text{active}\big{)}>0.$ (14)
A.2 Theory for complex question performance gain
A.2.1 Preliminaries
@@
$\mathcal{L}\left[p(y|x)\right]=$ $\sum_{x,y}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)}-\beta\sum_{x,y,c}p(x)p(y|x)p(c|x)\ln\frac{p(c|y)}{p(c)}$ (20)
$+\sum_{x}\lambda(x)\left(\sum_{y}p(y|x)-1\right).$
@@
$\frac{\partial}{\partial p(y|x)}\Bigl{(}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)}\Bigr{)}=p(x)\left[\ln\frac{p(y|x)}{p(y)}+1\right],$ (21)
@@
$\frac{\partial}{\partial p(y|x)}\Bigl{(}-\beta p(x)p(y|x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\Bigr{)}=-\beta p(x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}.$ (22)
$p(x)\left[\ln\frac{p(y|x)}{p(y)}+1\right]-\beta p(x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}+\lambda(x)=0.$ (23)
@@
A.2.2 proof of theorem 3.1
@@
From lemma A.1, by constructing the Lagrangian
$\mathcal{L}\bigl{[}p(e_{p}|e_{c})\bigr{]}=\sum_{e_{c},e_{p}}p(e_{c})p(e_{p}|e_{c})\ln\frac{p(e_{p}|e_{c})}{p(e_{p})}-\beta I(e_{p};C)+\sum_{e_{c}}\lambda(e_{c})\left(\sum_{e_{p}}p(e_{p}|e_{c})-1\right),$ (29)
@@
Table 5: Dataset specifications for SCORE framework
Category Details
Total Volume • 5,000 total episodes • 1,000 complete stories
Genre Distribution • Sci-Fi: 250 stories (25%) • Drama: 250 stories (25%) • Fantasy: 250 stories (25%) • Comedy: 250 stories (25%)
Structural Features • 10–15 eps/story ($\mu=12$) • 800–5k tokens/episode • 2k tokens avg./story ($\sigma=450$)
Generation • GPT-3.5 (50%) + GPT-4 (50%) • 15 prompts/genre • 3 annotators ($\kappa=0.78$)
Complexity • 30% Multi-char(>4) • 20% Non-linear • 50% Symbolic objects
Benchmarks • NarrativeQA: 1,567 stories • BookCorpus: 11k novels • WP-STORIES benchmark
Standardization • Genres: SF (20%), D (30%), F (30%), C (20%) • Length: 1.5k–2.5k tokens

### v2

@@ 5 Conclusion
We introduced SCORE, a novel LLM-based evaluation framework aimed at improving the coherence and emotional consistency of AI-generated stories. By incorporating Retrieval-Augmented Generation (RAG), TF-IDF, cosine similarity, and sentiment analysis, our method retrieves contextually relevant information, enabling more accurate and detailed assessments.
Experimental results demonstrate that SCORE outperforms traditional methods by effectively identifying continuity errors, maintaining narrative coherence, and providing valuable insights for episodic storytelling. However, limitations such as dependency on retrieval accuracy and computational demands present opportunities for future improvement.
In future work, we aim to optimize retrieval accuracy, refine emotional analysis, and expand the framework’s applicability to a broader range of LLM-generated stories to further validate its adaptability and generalizability.
Appendix A Implementation Settings
Table 3 to 6 display the results from both using the GPT API and directly utilizing the GPT playground to analyze the article. In the GPT playground, the full episode script is uploaded, and the prompt "Please evaluate episode n" is used to generate baseline results. When calling the GPT API, we upload the full episode and processed files, employ the OpenAI API combined with our framework to answer user queries, focusing primarily on evaluating the coherence of each episode. As noted earlier, the evaluation process has been optimized, resulting in a 90% improvement in analysis accuracy.
Appendix B Theoretical Analysis
B.1 Proof of the Continuity Analysis of Critical Item State Correction
@@
$\begin{split}&P\bigl{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1},\dots,S(t_{0})=s_{0}\bigr{)}\\ &=P\bigl{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1}\bigr{)}\,.\end{split}$ (1)
@@
Thus, no transition from destroyed or lost back to active or any non-terminal state is allowed.
@@
$\begin{split}\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}&=\\ -\ln\Bigl{(}P\bigl{(}S_{k+1}=\text{active}\mid S_{k}=\text{destroyed}\bigr{)}\Bigr{)}&>0\,.\end{split}$ (7)
@@
$\begin{split}E&=\sum_{j\neq k}\Delta E\bigl{(}S_{j},S_{j+1}\bigr{)}+\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}\\ &>0\,.\end{split}$ (8)
B.2 Theory for complex question performance gain
B.2.1 Preliminaries
@@
$\begin{split}\mathcal{L}\bigl{[}p(y\mid x)\bigr{]}&=\sum_{x,y}p(x)\,p(y\mid x)\,\ln\frac{p(y\mid x)}{p(y)}\\ &\quad-\beta\sum_{x,y,c}p(x)\,p(y\mid x)\,p(c\mid x)\,\ln\frac{p(c\mid y)}{p(c)}\\ &\quad+\sum_{x}\lambda(x)\Bigl{(}\sum_{y}p(y\mid x)-1\Bigr{)}\,.\end{split}$ (14)
@@
$\begin{split}&\frac{\partial}{\partial p(y\mid x)}\Bigl{(}p(x)\,p(y\mid x)\,\ln\tfrac{p(y\mid x)}{p(y)}\Bigr{)}\\ &\quad=p(x)\Bigl{[}\ln\tfrac{p(y\mid x)}{p(y)}+1\Bigr{]}\,.\end{split}$ (15)
@@
$\begin{split}&\frac{\partial}{\partial p(y\mid x)}\Bigl{(}-\beta\,p(x)\,p(y\mid x)\,\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\Bigr{)}\\ &\quad=-\beta\,p(x)\,\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\,.\end{split}$ (16)
$\begin{split}&p(x)\Bigl{[}\ln\tfrac{p(y\mid x)}{p(y)}+1\Bigr{]}-\beta\,p(x)\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\\ &\quad+\lambda(x)=0\,.\end{split}$ (17)
@@
Theorem B.2 (Optimal Retrieval Distribution via the Information Bottleneck Principle).
and the optimal retrieval probability is
$p(e_{p}|e_{c})=\frac{K(e_{c},e_{p})}{\sum_{e\in\mathcal{E}}K(e_{c},e)},$ (23)
The details and underlying proof are provided in the section B.2.
B.2.2 proof of theorem B.2
@@
From lemma B.1, by constructing the Lagrangian
$\begin{split}\mathcal{L}\bigl{[}p(e_{p}\mid e_{c})\bigr{]}&=\sum_{e_{c},e_{p}}p(e_{c})\,p(e_{p}\mid e_{c})\,\ln\frac{p(e_{p}\mid e_{c})}{p(e_{p})}\\ &\quad-\beta\,I(e_{p};C)\\ &\quad+\sum_{e_{c}}\lambda(e_{c})\Bigl{(}\sum_{e_{p}}p(e_{p}\mid e_{c})-1\Bigr{)}\,.\end{split}$ (25)
@@
Table 2: Evaluation Metrics Results
Evaluation Method BLEU-4